# MatFormer: Nested Transformer for Elastic Inference

### v1

@@ 4.2.1 Image Classification
We then explore using MatFormer at different training stages with a $2\times 2$ grid of pretraining-finetuning pairs (Table 7 in Appendix E.1) and find that using a MatFormer during pretraining helps bring more accurate and flexible encoders for downstream use. Further, finetuning using MatFormer enhances elastic deployment depending on the constraints at hand through Mix’n’Match.
During deployment, the universal MatViT model can be stored in memory and depending on the compute constraints be used to extract an adaptable smaller model to maximize accuracy with the available resources at that moment. Currently, we find the Mix’n’Match models on the accuracy-compute curve through a quick inference on the validation set. While relatively scalable, this points to the need for optimal budget allocation across layers in neural networks (Kusupati et al., 2020).
@@ 4.2.2 Adaptive Image Retrieval
The goal of image retrieval is to find semantically similar images – e.g. images from the same class – using representations obtained from a pretrained encoder (Chen et al., 2022). Standard approach is to encode the database images as well as query image with same encoder and run nearest neighbor retrieval for the query embedding. While we can embed database images with an expensive encoder, the query encoder generally has to be real-time. Furthermore, the setting of query encoding might be varied, e.g., on-device vs. cloud processing, varying query load and query complexity. Current solutions have to stick to a fixed encoder thus compromising on accuracy or cost for various settings.
@@
We evaluate both ViT and MatViT encoders on ImageNet-1K for image retrieval. We compute 1-nearest neighbor (NN) accuracy using the representation vector of the [CLS] token (also see Appendix E.2). Figure 5 shows that submodels extracted from MatViT can approximately preserve distances and provide significantly more flexibility. For example, with a loss of $<0.5\%$ accuracy, MatViT-L/16 can reduce compute cost by $40\%$. To our knowledge, this is the first result of its kind and opens up a wide variety of adaptive inference strategies for large-scale semantic search.
5 Conclusions
In this work we presented MatFormer, a natively elastic Transformer architecture that allows training a single universal model which can be used to extract hundreds of smaller accurate submodels at zero additional cost. We find that the MatFormer Language Model (MatLM) matches the perplexity & 1-shot accuracy of independently trained models. In fact, MatLM demonstrates an interesting loss-vs-compute scaling curve that is nearly independent of trained granularity indicating robust generalization to extremely large models as well. Finally, MatFormer submodels enable diverse inference time speedups like faster autoregressive generation with speculative decoding and elastic query encoders for adaptive dense retrieval across modalities.
@@ Acknowledgments
We are grateful to Aishwarya P S, Yashas B.L. Samaga, Varun Yerram, Lovish Madaan, Anurag Arnab for assistance in setting up training pipelines, Mattew Wallingford, Praneeth Netrapalli, Orhan Firat, Tom Duerig, Luke Zettlemoyer, Manish Gupta, Rahul Sukthankar and Jeff Dean for helpful discussions, support and feedback. We also acknowledge the computing resources and support from HYAK at the University of Washington, FAS RC at Harvard University, and Kempner Institute for the early-stage exploration of this project.
Appendix A Implementation Details
A.1 Architecture and Training
For our experiments, we train a range of MatLMs varying from size 78M to 2.6B for 10B-160B tokens – we scale model size equally with the number of training tokens (Hoffmann et al., 2022). For each MatLM granularity, we also train a corresponding baseline vanilla Transformer model. That is, for each model size we train Baseline-XL, L, M, S with $d_{ff}=4*d_{model},2*d_{model},d_{model},d_{model}/2$. All models have 16 layers, 16 attention heads, and a $d_{model}:d_{ff}$ ratio of $1:4$. We train a 256k vocabulary using the SentencePiece library (Kudo & Richardson, 2018), use a maximum context length of 1024 tokens, and a batch size of 1M tokens. We pretrained the 2.6B on 256 v3 TPU chips. We provide further details on these models in Table 3. For further details on training data, we point the reader to (Thoppilan et al., 2022).
@@ A.1 Architecture and Training
1.3B (1B / 927M / 860M) 805M (537M/403M/335M) 537M 2048 120B
2.6B (2B / 1.7B / 1.54B) 1.8B (1.2B/0.9B/0.7B) 1.2B 3072 160B
A.2 Downstream Evaluation
We evaluate all the LM models trained on set of 26 English tasks similar to (Brown et al., 2020; Du et al., 2022; Chowdhery et al., 2022; Anil et al., 2023), including:
@@ A.2 Downstream Evaluation
    Open-Domain Closed-Book Question Answering tasks: TriviaQA (Joshi et al., 2017), Natural Questions (Kwiatkowski et al., 2019), and WebQuestions (Berant et al., 2013).
@@
    Cloze and completion tasks: LAMBADA (Paperno et al., 2016), HellaSwag (Zellers et al., 2019), and StoryCloze (Mostafazadeh et al., 2016).
@@
    Winograd-style tasks: Winograd (Levesque et al., 2012) and WinoGrande (Sakaguchi et al., 2019).
@@
    Reading comprehension: SQuAD v2 (Rajpurkar et al., 2018) and RACE (Lai et al., 2017).
@@
    Common sense reasoning: PIQA (Bisk et al., 2019), ARC (Clark et al., 2018), and OpenBookQA (Mihaylov et al., 2018).
@@
    Natural language inference: Adversarial NLI (Nie et al., 2020).
Among all the downstream datasets, we classify LAMBADA, Natural Questions, SQuAD v2, WebQuestions, and TriviaQA under “GEN” tasks as these require generating a few tokens, and the remaining tasks under “RANK” tasks as they consist of choosing an option among the choices given along with the input. For all the granularities corresponding to each model, we present evaluation numbers along with development set log perplexity loss on all the 26 tasks in Tables 9, 10, 11, 12, 13, 14 and 15. We also perform evaluation on 2.6B Mix’n’Match models and provide it in Table 16.
Appendix B Training and Inference Costs
@@
We currently make minimal changes and optimizations to the training scripts of vanilla Transformer architecture. In other words, we use the same code for both Baselime and MatFormer, except using different sized splices of FFN block for each forward pass. Note that this implementation is suboptimal, as it involves added communication costs of FFN weight matrices when using model parallel training (discussed in more details in Section B.1). Though using a suboptimal implementation, we achieve the wall-clock time for MatLM training $\sim 15\%$ less to sum of wall-clock times to train all the $4$ granulatities baseline counterparts. We give exact FLOP count, wall-clock time, and forward pass time (inference cost) of each baseline and MatLM 2.6B model (or its corresponding smaller granularities) in Table 4. During serving, we observe the 2.6B model FFN latency to attention latency ratio $=56:44$. We emphasize that though we trained one MatFormer and compare its training time with Baselines combined, we get many more model than the 4 model granularities we explicitly trained for.
Table 4: 2.6B MatLM and Baseline training time per step, GFLOPs per step, and forward pass latencies. Each model is trained on 256 v3 TPU chips. Note that MatLM Fwd pass latency for any granularity will be same as corresponding Baseline granularity latency.
Model Time (s) / step GFLOPs / step Fwd pass latency (s)
MatLM 2.326 470841 -
Baseline-XL 0.728 186884 0.234
Baseline-L 0.670 147317 0.215
Baseline-M 0.652 125517 0.198
Baseline-S 0.630 117556 0.190
B.1 Improving MatFormer Training Efficiency
While MatFormer training uses asymptotically 2×2$\times$2 × FLOPs compared to a regular Transformer, optimizations are necessary to also realize a 2×2$\times$2 × runtime training performance. We discuss a few strategies here, leaving exact experimental testing to future work.
Delayed gradient synchronization via local accumulation. Since multiple forward and backward passes are made for each mini-batch in common implementations of data parallelism, this induces a gradient synchronization across all device for each backward pass with additional gradient accumulation. As such, for MatFormers a minimum of 2×2$\times$2 × the parameters worth of gradients are exchanged for the MLP layers, thus increasing the communication overhead. Additionally, for some frameworks, such as PyTorch, gradients of the full-weight matrix size need to be exchanged, leading to 4×4$\times$4 × more communication for our default experimental setup. A more efficient way to communicate gradients is to keep a local gradient accumulation buffer, which is used to accumulate all gradient from all subnetworks into the main, full-sized weight gradient. After all forward-backward passes have been completed, synchronization of gradients – with additional overall of computation and communication – can ensue. This saves 2×2$\times$2 × communication overhead, reducing communication overhead to the same cost as a regular Transformer.
Fused MatFormer kernels. Depending on the accelerator (GPU/TPU), the smallest MatFormer forward and backward pass can be inefficient in that the matrices are too small to fully utilize the accelerator. To improve utilization at the cost of additional memory for activations, it is possible to run the following computational fusion strategy for MatFormer computation: (a) duplicate mini-batch 4×4$\times$4 ×, (b) do the forward/backward pass for each layer for all MatFormer stages at the same time, (c) in doing so, load the tile for the weight matrix once, and reuse it for all relevant MatFormer stages. This strategy is similar to tiling strategies in FlashAttention (Dao et al., 2022) or convolution (Krizhevsky, 2009) which increase the arithmetic intensity for small weights by reusing of matrix multiplication tiles written to SRAM.
B.2 Speculative Decoding Attention Sharing
@@
An additional benefit of MatLM is that the attention cache is shared between the draft and verifier model. When the XL model verifies S model’s draft, it overwrites the attention cache with its richer latent representation compared to the one generated by the drafter model. Note that 1) this does not involve extra computation since MatLM has a single universal model including both draft and verifier model; 2) attention sharing isn’t possible in Baseline since they are not explicitly trained together. Hence, latent representation of one model is quite meaningless to the other model. Thus, attention sharing gives further improvement over vanilla speculative decoding as shown in Table 1.
Appendix C Scaling Laws for Language Decoders

### v2

@@ 4.2.1 Image Classification
We then explore using MatFormer at different training stages with a $2\times 2$ grid of pretraining-finetuning pairs (Table LABEL:tab:vit-2x2 in Appendix G.1) and find that using a MatFormer during pretraining helps bring more accurate and flexible encoders for downstream use. Further, finetuning using MatFormer enhances elastic deployment depending on the constraints at hand through Mix’n’Match.
@@
During deployment, the universal MatViT model can be stored in memory and depending on the compute constraints be used to extract an adaptable smaller model to maximize accuracy with the available resources at that moment. Currently, we find the Mix’n’Match models on the accuracy-compute curve through a quick inference on the validation set. While relatively scalable, this points to the need for optimal budget allocation across layers in neural networks [33].
@@ 4.2.2 Adaptive Image Retrieval
The goal of image retrieval is to find semantically similar images – e.g. images from the same class – using representations obtained from a pretrained encoder [13]. Standard approach is to encode the database images as well as query image with same encoder and run nearest neighbor retrieval for the query embedding. While we can embed database images with an expensive encoder, the query encoder generally has to be real-time. Furthermore, the setting of query encoding might be varied, e.g., on-device vs. cloud processing, varying query load and query complexity. Current solutions have a fixed encoder thus compromising on accuracy or cost for various settings.
@@
We evaluate both ViT and MatViT encoders on ImageNet-1K for image retrieval. We compute 1-nearest neighbor (NN) accuracy using the representation vector of the [CLS] token (also see Appendix G.2). Figure 5 shows that submodels extracted from MatViT can approximately preserve distances and provide significantly more flexibility. For example, with a loss of $<0.5\%$ accuracy, MatViT-L/16 can reduce compute cost by $40\%$. This corresponds to the 175M Mix’n’Match parameter model in Fig 5(b), which is 40% smaller than the 300M XL model, and has $<0.5\%$ accuracy drop. To our knowledge, this is the first result of its kind and opens up a wide variety of adaptive inference strategies for large-scale semantic search.
5 Conclusion
In this work we presented MatFormer, a natively elastic Transformer architecture that allows training a single universal model which can be used to extract hundreds of smaller accurate submodels at zero additional cost at deployment time. We find that the MatFormer Language Model (MatLM) matches the perplexity & 1-shot accuracy of independently trained models. In fact, MatLM demonstrates an interesting loss-vs-compute scaling curve that is nearly independent of trained granularity indicating robust generalization to extremely large models as well. Finally, MatFormer submodels enable diverse inference time speedups like faster autoregressive generation with speculative decoding and elastic query encoders for adaptive dense retrieval across modalities. We believe dynamically routing these models to change inference latency [32, 40, 20], and developing the hardware optimizations required is a promising area for future work.
@@ Acknowledgments
We are grateful to Aishwarya P S, Yashas B.L. Samaga, Varun Yerram, Lovish Madaan, Anurag Arnab for assistance in setting up training pipelines, Matthew Wallingford, Praneeth Netrapalli, Orhan Firat, Rohan Anil, Tom Duerig, Luke Zettlemoyer, Manish Gupta, Rahul Sukthankar and Jeff Dean for helpful discussions, support and feedback.
We also acknowledge the computing resources and support from HYAK at the University of Washington, FAS RC at Harvard University, Kempner Institute and a GCP credit award for the early-stage exploration of this project. Ali Farhadi acknowledges funding from the NSF awards IIS 1652052, IIS 1703166, DARPA N66001-19-2-4031, DARPA W911NF-15-1-0543, and gifts from Allen Institute for Artificial Intelligence and Google. Sham Kakade acknowledges funding from the Office of Naval Research under award N00014-22-1-2377. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. Yulia Tsvetkov acknowledges support from the National Science Foundation under CAREER Grant No. IIS2142739, NSF grants No. IIS2125201 and IIS2203097, and gift funding from Google, MSR, and OpenAI. Hannaneh Hajishirzi acknowledges funding through a gift from Allen Institute for Artificial Intelligence.
Appendix A Broader Impact Statement
The elasticity of MatFormer enables its use in diverse deployment scenarios. This lowers the barrier for practitioners to use foundation models tailored to their deployment scenarios. Training foundation models remains expensive, with the largest models we discuss trained on 256 TPU-v4 cores for 3 days. Moreover, while we report validation loss and downstream evaluation scores on a variety of tasks, we acknowledge the possibility that MatFormer can have adverse effects on bias [26] or underrepresented domains/languages [1].
Appendix B Implementation Details
B.1 Architecture and Training
For our experiments, we train a range of MatLMs varying from size 78M to 850M for 10B-80B tokens – we scale model size equally with the number of training tokens [25]. For each MatLM granularity, we also train a corresponding baseline vanilla Transformer model. That is, for each model size we train Baseline-XL, L, M, S with $d_{ff}=4*d_{model},2*d_{model},d_{model},d_{model}/2$. All models have 16 layers, 16 attention heads, and a $d_{model}:d_{ff}$ ratio of $1:4$. We train a 256k vocabulary using the SentencePiece library [31], use a maximum context length of 1024 tokens, and a batch size of 1M tokens. We pretrained the 850M models on 256 v3 TPU chips. We provide further details on these models in Table 4. For further details on training setup, we point the reader to [58].
@@ B.1 Architecture and Training
B.2 Downstream Evaluation
We evaluate all the LM models trained on set of 25 English tasks similar to [8, 22, 14, 3], including:
@@ B.2 Downstream Evaluation
    Open-Domain Closed-Book Question Answering tasks: TriviaQA [28], Natural Questions [35], and WebQuestions [4].
@@
    Cloze and completion tasks: LAMBADA [46], HellaSwag [67], and StoryCloze [43].
@@
    Winograd-style tasks: Winograd [38] and WinoGrande [51].
@@
    Reading comprehension: RACE [37].
@@
    Common sense reasoning: PIQA [6], ARC [15], and OpenBookQA [42].
@@
    Natural language inference: Adversarial NLI [44].
For all the granularities corresponding to each model, we present evaluation numbers along with development set log perplexity loss on all the 25 tasks in Tables 9, 10, 11, 12 and 13.
B.3 Baseline Implementation Details
Independently trained Baseline: Baselines are trained from scratch independently, where each granularity in a given model size uses X tokens mentioned in Table 4. For example, for 850M model X = 80B tokens. Therefore, Baseline-{S, M, L, XL} process 4X tokens in total. Baselines trained from scratch result in the same number of models that are explicitly trained for. That is, there is no way to directly get a submodel of size in between the trained granularities without using additional methods like distillation, pruning, etc.
OFA: OFA [9] trains a big model from scratch, then freezes it and starts another run of a similar sized "universal model" where it samples random subnetworks and optimizes them with distillation and ground truth loss. The central idea of this work is sampling random submodels and optimizing them. Distillation is a orthogonal component which can be added to MatFormer as well. In order to use comparable compute and memory to the Baseline (Section 4.1), we modify the method to only optimize w.r.t. ground truth loss - as also done in the case of MatFormer. Overall, OFA uses 4X tokens, same as all the baseline granularities trained form scratch. In each step, OFA samples a random model, where each of its layer has {S, M, L, XL} granularity sampled randomly, and optimizes it.
To obtain submodels according to a given cost constraint, OFA proposes using NAS [48] to search for optimal architecture. Specifically, OFA samples a random set of (architecture, loss) pairs from the universal model. Then, it trains a predictor on this pair, taking architecture as input and predictiing the loss. Finally, it uses an evolutionary search based routine [48] to search for the optimal architecture within a specified constraint, using the predictor in its search routine. More details of NAS and comparison with MatFormer Mix’n’Match is presented in Section D.
DynaBERT: DynaBERT [27] has $g$ fixed granularities of submodels - similar to MatFormer. Similar to OFA, DynaBERT employs distillation loss as well. For a comparable analysis and to maintain compute/memory comparable to baseline, we only optimize DynaBERT wrt ground truth loss, similar to MatFormer.
The crucial difference between DynaBERT and MatFormer lies in its training methodology. In each step, DynaBERT takes 4 batches of data, optimizes each batch with each of the {S, M, L, XL} granularity and averages these losses. Similar to Matformer and OFA, it processes a total of 4X tokens. Though, since it optimizes on 4 batches of data in a single step, it runs for one fourth the steps of MatFormer or OFA, and same number of steps as the baseline.
An important distinction between contribution of DynaBERT work and MatFormer is the search strategy. In MatFormer, we introduced the Mix’n’Match heuristic, which calculates an optimal submodel with a cost constraint without any overhead. In contrast, DynaBERT doesn’t introduce any search strategy and discusses only using explicitly trained granularities as submodels.
B.4 Comparison with Baselines
Here we do a more detailed discussion of comparison between MatFormer and the baselines used in Section 4.
1. MatFormer vs trained from scratch Baseline: Let Baseline-{S, M, L, XL} combined process 4X tokens, where each granularity is trained on X tokens independently. Since MatFormer maintains just one model, we train it for 4X tokens, in each step sampling either of MatLM-{S, M, L, XL}. As a result, the total compute and peak memory is still the same as Baseline since we effectively train MatLM-{S, M, L, XL} for X tokens each – matching Baseline training configuration. But since we maintain a single model, the shared parameters get trained for up to 4X tokens for free. For example, each time we train XL/L/M granularity, since S granularity is contained in them, it also receives gradients updates. Therefore, the MatLM-S effectively processes 4X tokens, MatLM-M processes 3X tokens, MatLM-L processes 2X tokens, and MatLM-XL processes X tokens. We can also see this translating to Validation loss (Figure 2(a)) – the smaller granularities outperform Baseline by a large margin. Whereas the larger granularity match in performance.
2. MatFormer vs DynaBERT: DynaBERT, similar to MatFormer, trains $g$ nested subnetworks, but differs critically in its training methodology. It performs joint optimization - simultaneously processing $4$ batches of data - $1$ batch through each {S, M, L, XL} granularity. The final loss being optimized is the average of each granularity’s loss. Conversely, MatFormer samples and optimizes individual granularities in separate steps, resulting in four times more gradient updates for the same amount of data and compute. This difference in training methodology translates to a substantial performance gap: DynaBERT exhibits a significant 0.01 log perplexity gap compared to MatFormer on the 850M model. Even with a 15% increase in compute, DynaBERT still falls short of MatFormer’s performance. Moreover, unlike MatFormer, which offers flexibility in submodel size selection through its Mix’n’Match approach, DynaBERT lacks a dedicated sampling strategy.
3. MatFormer vs OFA: OFA maintains one universal model similar to MatFormer, but randomly samples subnetworks such that each layer is one of {S, M, L, XL}. Similar to MatFormer and DynaBERT, it takes advantage of processing 4X data by a single model. But due to sampling, not enough small or large models (close to S and XL granularity model) are sampled, leading to inferior performance in these regions. That is, sampling models ranging from S model size to XL model size, most of the sampled models are close to M-L model sizes. This can also be seen in the Loss curve, where we see a bell like curve having much higher loss close to S and XL model sizes, and matching MatLM’s performance in between. Another disadvantage of OFA is since random model are sampled, once can’t employ simple Mix’n’Match technique like in matFormer. Rather a more complicated NAS strategy is needed to find optimal submodels. To train a NAS when searching for such large models is costly and erroneous, as we discuss in Appendix D.2.
Appendix C Training and Inference Costs
@@
We currently make minimal changes and optimizations to the training scripts of vanilla Transformer architecture. In other words, we use the same implementation for both Baselime and MatFormer, except using different sized splices of FFN block for each forward pass. The wall-clock time for MatLM training is the same cost as training all the $4$ granulatities baseline counterparts. During serving, we observe the 850M model FFN latency to attention latency ratio $=56:44$. We note that this FFN:MHA latency ratio depends highly on scale and sequence length. More specifically, for a given sequence length FFN latency dominates the overall latency at scale, while the attention heads’ cost increases with sequence length. We refer the reader to Kim et al. [30] for a more extensive illustration of this. We emphasize that though we trained one MatFormer and compare its training time with Baselines combined, we can extract many more models than the 4 model granularities we explicitly trained.
C.1 Speculative Decoding Attention Sharing
@@
An additional benefit of MatLM is that the attention cache is shared between the draft and verifier model. When the XL model verifies S model’s draft, it overwrites the attention cache with its richer latent representation compared to the one generated by the drafter model. Note that 1) this does not involve extra computation since MatLM has a single universal model including both draft and verifier model; 2) attention sharing isn’t possible in the Baseline since they are not explicitly trained together. Hence, latent representation of one model is quite meaningless to the other model. Thus, attention sharing gives further improvement over vanilla speculative decoding as shown in LABEL:tab:spec.
Figure 6: We search for optimal architectrure configuration using evolutionary search [48] and with Mix’n’Match. We find that Mix’n’Match yields architectures lying on pareto-optimal curve, and are at least as good as those found using the evolutionary search.
Table 5: On 850M MatFormer model, while running NAS we observe it prefers balanced granularities across layers rather than skewed. On a few parameter constraints we list the MnM heuristic configuration, and configuration predicted with NAS.
FFN Params budget Mix’n’Match configuration & loss NAS predicted configuration & loss
226M [M,M,M,M,M,M,M,M,M,M,M,M,M,M,M,M]; 2.931 [S,M,M,M,M,M,M,S,M,M,M,M,M,M,M,L]; 2.944
245M [M,M,M,M,M,M,M,M,M,M,M,M,L,L,L,L]; 2.926 [M,M,M,M,M,L,M,M,M,L,M,L,M,M,M,L]; 2.932
278M [M,M,M,M,M,L,L,L,L,L,L,L,L,L,L,L]; 2.9 [M,L,L,L,L,L,L,L,L,L,M,L,M,M,M,L]; 2.91
377M [L,L,L,L,XL,XL,XL,XL,XL,XL,XL,XL,XL,XL,XL,XL]; 2.865 [L,L,XL,L,XL,L,XL,L,XL,XL,XL,L,XL,L,L,XL]; 2.874
Appendix D Search Techniques
D.1 Mix’n’Match
For a given compute or parameter budget multiple submodels may meet the constraint. A common strategy could involve employing Neural Architecture Search (NAS) to identify the optimal architecture within this subset; however, this approach can be prohibitively expensive. Instead, we propose a simpler Mixn’Match heuristic, which identifies optimal submodels that lie on the Pareto-optimal accuracy-vs-parameters curve.
The Mix’n’Match heuristic recommends selecting layer granularities with minimal changes across layers, ensuring that the granularity of the $j^{th}$ layer is at least that of the $i^{th}$ layer for $j>i$. This means that a gradual increase in granularity with the "least slope" empirically yields the best performance among all tested mix-and-match configurations. Our heuristic is underpinned by the training methodology, where each sampled granularity configuration maintains consistent layer granularity across the model. Consequently, the model adapts best to configurations where layer granularities are either uniform or display minimal variation. For instance, a configuration of [M, M, M, M, L, L, L] across layers is more effective than [S, S, S, S, S, XL, XL] despite having similar number of parameters, as it maintains a more balanced distribution of granularity as opposed to a skewed one.
This intuition is also supported by results from an evolutionary search-based Neural Architecture Search (NAS) method [48], similar to that employed in [9]. Our NAS experiments, few of the runs we present in Table 5, indicated a preference for balanced configurations under various constraints, aligning with the findings from our Mix’n’Match heuristic.
We explored various balanced configurations, including:
  1. 1.
    Increasing-Decreasing: Granularity increases until the midpoint and then decreases, e.g., [M, M, L, L, L, M, M].
  2. 2.
    Decreasing-Increasing: The reverse of the previous heuristic, where granularity decreases then increases.
  3. 3.
    Increasing: A non-decreasing sequence of layer granularities, such as [M, M, M, M, L, L, L].
  4. 4.
    Decreasing: The opposite of the Increasing heuristic.
Among these configurations, the Increasing heuristic consistently outperformed the others. Therefore, we recommend adopting a strategy of selecting increasing granularities across layers with the minimum slope for effective model performance.
On the 850M MatLM, we compare architectures predicted by Mix’n’match and the NAS algorithm, in Figure 6. We find that our heuristic works at least as well as NAS, resulting in models that lie on pareto-optimal curve. Note that Mix’n’match requires no additional training of a module or overhead in calculating optimal submodel within a cost constraint.
D.2 NAS
Neural Architecture Search [68, 48] is a natural method to search an optimal network among possible architectures within a constraint budget. In fact, OFA [9] uses an evolutionary search based technique [48] along with a architecture loss predictor (Appendix B.3) in its search routine. While on smaller scale models, it might be easy to deploy NAS method similar to what OFA does, on larger scale it becomes prohibitively costly and erroneous.
To get loss corresponding to an architecture, one needs to run on a sufficiently large held out set to get an average value. For example, in our runs we used a held out set of around 200M tokens. On large scale, it becomes costly to repeatedly run large models on this held out set to collect loss values. For example, to collect  16k (architecture, loss) pairs, we used 256 TPUv-4 chips for around 2 days.
Once we have the dataset, we split it into 60%-40% train-eval split. After training the architecture loss predictor, we tested out its performance on the held out set and noticed $\sim 5e-5$ mean squared error.
As we move to large models regime, the performance gap between different sized models diminish. For example, the gap between 78M parameter Baseline-S and Baseline-XL, which are of size 71M and 78M, is 4.01 vs 3.83. Which is a gap of 0.18 loss for just 7M parameter difference. But in case of 850M model, Baseline-S and Baseline-XL has size 582M vs 850M , the corresponding loss is 3.017 vs 2.84. Thus, a difference of 0.177 for 268M parameter difference. This trend scales to larger models, and the gap reduces between various model sizes. So, training a architecture accuracy predictor becomes increasingly challenging and erroneous, since it has to learn differentiating architectures with very small differences in their loss values.
In our experiment for OFA, we take the best of the model configs predicted by NAS and the model config among the sampled points for a corresponding budget.
Appendix E Scaling Laws for Language Decoders